# Survey on Publicly Available Sinhala Natural Language Processing Tools and Research

### v22

@@ 3.18.2 Handwritten Text
Fernando et al. [407] claim to have created a database for handwriting recognition research in the Sinhala language and further claims that the data set is available at the National Science Foundation (NSF) of Sri Lanka. However, the paper provides no URLs and we were not able to find the data set on the NSF website either. The work by Karunanayaka et al. [408] is focused on noise reduction and skew correction of Sinhala handwritten words. A genetic algorithm-based approach for non-cursive Sinhala handwritten script recognition was proposed by Jayasekara and Udawatta [409]Nilaweera et al. [410] compare projection and wavelet-based techniques for recognizing handwritten Sinhala script.  Silva and Kariyawasam [411] worked on segmenting Sinhala handwritten characters with special focus on handling diacritic vowels. A comparative study of few available Sinhala handwriting recognition methods was done by Silva et al. [412]Silva et al. [413] uses contour tracing for isolated characters in handwritten Sinhala text. A Sinhala handwriting OCR system which utilizes zone-based feature extraction has been proposed by Dharmapala et al. [414]. The study by Walawage and Ranathunga [415] and its follow up study by Walawage [416] specifically focus on segmentation of overlapping and touching Sinhala handwritten characters. Silva and Jayasundere [417] focused on recognizing character modifiers in Sinhala handwriting. The similarly named studies by Mariyathas et al. [418] and Wasalthilake and Kartheeswaran [419], both utilize CNN to recognize Sinhala handwriting; as does the study by Weerasinghe [420]Ifhaam and Jayalal [421] used genetic algorithms to recognize Sinhala handwritten postal addresses for postal sorting. A segmentation-based approach that utilizes an n-gram model to recognize and validate Sinhala words written on touch screens was proposed by Mahesh and Priyankara [422]. They used a CNN classifier and were able to classify 19 different Sinhala characters. The study by Rowel et al. [423] frames their work as an E-Learning platform for hearing impaired children. However, their research does not contain any work done towards Sinhala sign language to be included in Section 3.23. What they do have is an OCR system that they claim to recognize letters and digits. Even there, we are only given an example of a recognised digit. Whether or not their system can recognize Sinhala letters is not explicitly discussed. As part of their child [sic] cognitive ability assessment model, Kahawanugoda et al. [351], proposed a Sinhala handwriting letter recognition system. The study by Withana and Rupasinghe [424] used Sinhala handwritten text classification to detect Dyslexia and Dysgraphia. The study by eka [425] has used CNN to recognize Sinhala handwritten text.
@@ 3.18.3 Ancient Text
Summarizing optically recognized old Sinhala text for the purpose of archival search and preservation was explored by Rathnasena et al. [426]. The work of Peiris [427] also focused on OCR for ancient Sinhala inscriptions. Building upon the architecture proposed by Ruwanmini et al. [428], a neural network-based method for recognizing ancient Sinhala inscriptions was proposed by Karunarathne et al. [429]. The study by Wickramarathna and Ranathunga [430] created a system to recognize Brahmi characters, correct errors, and generate Sinhala meanings. Heenkenda and Fernando [431] used Inception-v3 [432], VGG-19 [433], and ResNet-50 [434] to classify Sinhala inscriptions to historical time periods.
@@ 3.19 Translators
A meta-study on the viability of machine translators replacing English to Sinhala human translators was conducted by Dilshani and Senevirathna [435]. However, this study only involves 100 combined and complex English sentences translated to Sinhala by human translators as well as MT software. Given that reason and the fact that they seem to only used Google translate and Akura Sinhala dictionary app for comparison, the conclusions of this study may not be generalized. Another meta-study on the impact of pre-trained multilingual sequence-to-sequence models on low-resource language translation has been conducted by Lee et al. [436]; while they consider Sinhala as one of the examples, they do not go much into the specific impact due to the general nature of the paper. The meta-analysis by Ramadasa et al. [437] attempts to evaluate the goodness of the Google Sinhala-to-English translation by using the Google cloud API to translate Sinhala to English and then analysing the accuracy of the Sentiment Analysis task and the Named Entity Recognition task on the translated text. The meta-analysis by Das et al. [438] compared the results of translating English to 15 Indic languages including Sinhala using statistical translation methods. They used datasets from OPUS [123] for model building and utilized Flores-200 for fine-tuning. The NMT for Indic languages study conducted by Sheshadri et al. [439] discusses the Sinhala language translation in the abstract and conclusion but the paper itself focuses more on the languages spoken in India. Nevertheless, it puts Sinhala into a regional linguistic perspective. The meta-study by Bapna et al. [440] discusses the task of building clean, web-mined datasets for a number of languages including Sinhala for the task of machine translation. This discussion was continued by Jones et al. [441] who discussed the bilingual lexica (BILEXs) in the context of a number of languages including Sinhala.
@@
As mentioned in Section 3.1, the study by [96] raised questions on the quality of the existing Sinhala-English corpora. In a follow-up study by Ranathunga et al. [147], it was pointed out that using an automated ranking based on a similarity measure on web-mined corpora and using the resultant top samples can yield better translation models for English-Sinhala and Sinhala-Tamil. These models sometimes were better performing than a model trained on the full dataset. They further showed that using human labour to clean web-mined parallel corpora only gives marginal benefits over automated ranking and filtering. Thus they concluded that using expensive human labour for this task might not be efficient. Sen et al. [442] on the other hand attempted to improve the quality of the Sinhala-English parallel corpora using fuzzy string matching where they tried to match the English translation of the given Sinhala sentence to the English sentence in the dataset pair.
@@ 3.19.3 Singlish to Sinhala (Transliteration)
The XTREME-UP114114114https://github.com/google-research/xtreme-up data set created by Ruder et al. [134] contains a Sinhala data set for the transliteration task. The early work by Goonetilleke et al. [472] attempted Sinhala transliteration through the Latin alphabet. However, their work does not use the word transliteration and instead focuses on the predictive aspect. Priyadarshani et al. [473] used statistical machine learning for transliteration of names between Sinhala, Tamil, and English. A rule-based system on trigrams was proposed by Liwera and Ranathunga [474] for Singlish to Sinhala transliteration of social media text. A Singlish to Sinhala converter which uses an LSTM was proposed by De Silva [475]. A rule-based approach for the same was proposed by de Silva and Ahangama [476]. The study by Nanayakkara et al. [477] introduced an English-to-Sinhala transliteration system. Sumanathilaka et al. [478, 479] proposed a Trie [480] data structure-based algorithm for word suggestion in Sinhala transliterations. An extended analysis of the same work was presented in a later work [481]. Amarasekara et al. [482] proposed a rule-based method supported by N-gram analysis and a corpus dataset to transliterate Singlish tweets to Sinhala. The study by Rajapaksha et al. [483] claims to have created a translation system, however, what they have created is a transliteration system between Singlish and Sinhala. While they also claim to have trained ASR and TTS systems, the paper does not identify them by name or citation. The same is true for the data sets they have used to train the various modules in their pipeline. The studies by Kumaravithana et al. [484] and Jayawardhana et al. [485] claim to have built Sinhala-English translators but in actuality are just invoking Google API for their translation tasks.
The work by Kirov et al. [486] used the Dakshina [146] dataset and compared the accuracy of Singlish to Sinhala transliteration with vanilla LSTM, vanilla Transformers, mT5 [299], ByT5 [487], and a non-neural finite-state transducer (FST) based on work by Bisani and Ney [488]. They discuss in detail how the zero-width joiner (ZWJ, U+200D) character cause issues in this task and Sinhala rendering as a whole. Their code is available on GitHub115115115https://github.com/google-research/google-research/tree/master/context_aware_transliteration. Khiu et al. [489] fine-tuned mBART on Sinhala Government Corpus and Bible corpus to predict the NMT performance dependence on domain similarity.
@@ 3.19.5 Between Sinhala and Non-English Languages
Most of the cross-Sinhala and Tamil work has been done in the domain of machine translation. A neural machine translation for Sinhala and Tamil languages was initiated by Tennage et al. [493, 494]. Then they further enhanced it with transliteration and byte pair encoding [495] and used synthetic training data to handle the rare word problem [496]. This project produced Si-Ta [497] a machine translation system of Sinhala and Tamil official documents. In the statistical machine translation front, Farhath et al. [498] worked on integrating bilingual lists. The attempts by Weerasinghe [499] and Sripirakas et al. [500] were also focused on statistical machine translation while Jeyakaran and Weerasinghe [501] attempted a kernel regression method. A yet another attempt was made by Pushpananda et al. [502] which they later extended with some quality improvements [503]. An attempt at real-time direct translation between Sinhala and Tamil was done by Rajpirathap et al. [504]Dilshani et al. [505] have done a study on the linguistic divergence of Sinhala and Tamil languages with respect to machine translation. Mokanarangan [506] claims to have built a named entity translator between Sinhala and Tamil for official government documents. But this work is locked behind an institutional repository wall and thus is not accessible by other researchers. Arukgoda et al. [507] studied the possibility of using deep learning techniques to improve Sinhala-Tamil translation which they further improved later [508]Pramodya et al. [509] compared Transformers, Recurrent Neural Networks, and Statistical Machine Translation (SMT) in the context of Tamil to Sinhala machine translation. The work by Nissanka et al. [510] used monolingual word embedding to improve NMT between Sinhala and Tamil. Thillainathan et al. [511] uses pre-trained mBART [471] models for six directional translations between Sinhala, Tamil, and English. Yashothara and Uthayasanker [512] discussed the use of the Hierarchical Phrase-Based Model for Tamil to Sinhala and Sinhala to Tamil translations. Pramodya [513] presented a comparison of SMT [453] and NMT models for Sinhala-Tamil translation. For NMT they have used base transformer, turned transformer, and mT5 [299]. The work by Su et al. [514] which compared 8 Parameter-efficient fine-tuning (PEFT) methods on the Si-Ta language pair using the Government corpus [454] and NLLB [127] reports that Houlsby adapter [224], with a 33.34 SacreBLEU [515] score, to be the best PEFT method. However, they also note that the Pfeiffer adapter [226] runs the fastest at 52.59 hours with a reasonable score of 31.24. (Comparatively, the Houlsby adapter runs for 78.65 hours).
@@ 3.20 Spelling and Grammar
The open-source data driven approach proposed by Wasala et al. [522, 523] claims to be able to check and correct spelling errors in Sinhala. The approach by Jayalatharachchi et al. [524] attempts to obtain synergy between two algorithms for the same purpose. These efforts [522, 524] were then extended by Subhagya et al. [525]. A rule-based Sinhala spell checker named SinSpell based on Hunspell116116116http://hunspell.github.io/ was introduced by Liyanapathirana et al. [526]. They have also made the tool available117117117http://nlp-tools.uom.lk/sinspell/ online for use. The study by Sithamparanathan and Uthayasanker [527] extended the Generic Environment for context-aware spell correction to handle Sinhala and Tamil. Sonnadara et al. [120] created a benchmark data set for Sinhala spell correction along with a neural model. A multi agent-based spell checker, named LaSi Spell for Sinhala spell checking was introduced by Samarawickrama et al. [528]. The study by Udagedara et al. [529] specifically solved the problem of spell-checking Sri Lankan names and addresses. A system named Erroff was proposed by Sudesh et al. [530] to correct real-word errors in Sinhala text.
A model for detecting grammatical mistakes in Sinhala was developed by Pabasara and Jayalal [531]. They followed this up with a grammatical error detection and correction model [532]Gunasekara et al. [533] used annotation projection for semantic role labelling for Sinhala. A Sinhala grammar checker based on Hidden Markov models was developed by Fernando and Arudchelvam [534]. Widyaratna [535] used a sequence-to-sequence model with attention, which is generally used for translation tasks, to translate sentences with common grammatical errors to their corrected counterparts. The work by Jayasuriya et al. [536] used a rule-based approach along with Google Translation to correct grammar is Sinhala text. A rule-based system to convert Sinhala sentences from active voice to passive voice while correcting grammatical errors was proposed by Ilukkumbura and Rupasinghe [537]. The study by Goonawardena et al. [538] a rule-based system to spell-check Sinhala text as well as detect and correct grammatical errors. Navoda et al. [539] also claim to have created an automated tool to check Sinhala spelling and grammar.
@@ 3.21 Chat Bots
A simple Sinhala chat bot which utilizes a small knowledge base has been proposed by Hettige and Karunananda [70]. A study on the effect of word embeddings on a Sinhala chatbot was conducted by Gamage et al. [540] where they used, the fasttext model trained by Facebook [113, 114, 115], on a RASA118118118https://rasa.com/ [541] chat bot. A Sinhala chat bot for train information was proposed by Harshani [542]. Similarly, the tool proposed by Chandrasena et al. [543] serves as a chat bot-based recommendation system for Sri Lankan traditional dancers. The chat bot discussed by Kumanayake [544] has the very specific purpose of answering user inquiries about the degree programs at University of Ruhuna. Avishka et al. [545] used off the shelf RASA NLU Engine [541] and Microsoft Bot Network [546] to set up a generic Sinhala chat bot architecture. They demonstrated the effectiveness of their architecture by creating a food ordering chat bot. A web-based code-less chat bot development platform for Sinhala was proposed by Dissanayake et al. [547]. Further, they claimed that their system can handle OOV tokens as well as Sinhala-English code-switching. The work by Dasanayaka and Warnajith [548] used a deep learning Intent Mapping (IM) model to map consumer responses in their Sinhala chat bot framework. Rajapakshe et al. [362] proposed a Sinhala conversational interface for the purpose of scheduling medical appointments and giving medical advice. The chat bot component was in the back end while the front end contained speech recognition and voice synthesizer components.

### v23

@@ 3.18.2 Handwritten Text
Fernando et al. [423] claim to have created a database for handwriting recognition research in the Sinhala language and further claims that the data set is available at the National Science Foundation (NSF) of Sri Lanka. However, the paper provides no URLs and we were not able to find the data set on the NSF website either. The work by Karunanayaka et al. [424] is focused on noise reduction and skew correction of Sinhala handwritten words. A genetic algorithm-based approach for non-cursive Sinhala handwritten script recognition was proposed by Jayasekara and Udawatta [425]Nilaweera et al. [426] compare projection and wavelet-based techniques for recognizing handwritten Sinhala script.  Silva and Kariyawasam [427] worked on segmenting Sinhala handwritten characters with special focus on handling diacritic vowels. A comparative study of few available Sinhala handwriting recognition methods was done by Silva et al. [428]Silva et al. [429] uses contour tracing for isolated characters in handwritten Sinhala text. A Sinhala handwriting OCR system which utilizes zone-based feature extraction has been proposed by Dharmapala et al. [430]. The study by Walawage and Ranathunga [431] and its follow up study by Walawage [432] specifically focus on segmentation of overlapping and touching Sinhala handwritten characters. Silva and Jayasundere [433] focused on recognizing character modifiers in Sinhala handwriting. The similarly named studies by Mariyathas et al. [434] and Wasalthilake and Kartheeswaran [435], both utilize CNN to recognize Sinhala handwriting; as does the study by Weerasinghe [436]Ifhaam and Jayalal [437] used genetic algorithms to recognize Sinhala handwritten postal addresses for postal sorting. A segmentation-based approach that utilizes an n-gram model to recognize and validate Sinhala words written on touch screens was proposed by Mahesh and Priyankara [438]. They used a CNN classifier and were able to classify 19 different Sinhala characters. The study by Rowel et al. [439] frames their work as an E-Learning platform for hearing impaired children. However, their research does not contain any work done towards Sinhala sign language to be included in Section 3.23. What they do have is an OCR system that they claim to recognize letters and digits. Even there, we are only given an example of a recognised digit. Whether or not their system can recognize Sinhala letters is not explicitly discussed. As part of their child [sic] cognitive ability assessment model, Kahawanugoda et al. [363], proposed a Sinhala handwriting letter recognition system. The study by Withana and Rupasinghe [440] used Sinhala handwritten text classification to detect Dyslexia and Dysgraphia. The study by eka [441] has used CNN to recognize Sinhala handwritten text. Karunarathne et al. [442] used Gabor-initialized CNN (GCNN) on a dataset of 6000 handwritten Sinhala character images.
@@ 3.18.3 Ancient Text
Summarizing optically recognized old Sinhala text for the purpose of archival search and preservation was explored by Rathnasena et al. [443]. The work of Peiris [444] also focused on OCR for ancient Sinhala inscriptions. Building upon the architecture proposed by Ruwanmini et al. [445], a neural network-based method for recognizing ancient Sinhala inscriptions was proposed by Karunarathne et al. [446]. The study by Wickramarathna and Ranathunga [447] created a system to recognize Brahmi characters, correct errors, and generate Sinhala meanings. Heenkenda and Fernando [448] used Inception-v3 [449], VGG-19 [450], and ResNet-50 [451] to classify Sinhala inscriptions to historical time periods. The work by Gunasekara et al. [452] claims to have created a mobile app that is capable of recognizing and translating early Brahmi characters.
@@ 3.19 Translators
A meta-study on the viability of machine translators replacing English to Sinhala human translators was conducted by Dilshani and Senevirathna [453]. However, this study only involves 100 combined and complex English sentences translated to Sinhala by human translators as well as MT software. Given that reason and the fact that they seem to only used Google translate and Akura Sinhala dictionary app for comparison, the conclusions of this study may not be generalized. Another meta-study on the impact of pre-trained multilingual sequence-to-sequence models on low-resource language translation has been conducted by Lee et al. [454]; while they consider Sinhala as one of the examples, they do not go much into the specific impact due to the general nature of the paper. The meta-analysis by Ramadasa et al. [455] attempts to evaluate the goodness of the Google Sinhala-to-English translation by using the Google cloud API to translate Sinhala to English and then analysing the accuracy of the Sentiment Analysis task and the Named Entity Recognition task on the translated text. The meta-analysis by Das et al. [456] compared the results of translating English to 15 Indic languages including Sinhala using statistical translation methods. They used datasets from OPUS [126] for model building and utilized Flores-200 for fine-tuning. The NMT for Indic languages study conducted by Sheshadri et al. [457] discusses the Sinhala language translation in the abstract and conclusion but the paper itself focuses more on the languages spoken in India. Nevertheless, it puts Sinhala into a regional linguistic perspective. The meta-study by Bapna et al. [458] discusses the task of building clean, web-mined datasets for a number of languages including Sinhala for the task of machine translation. This discussion was continued by Jones et al. [459] who discussed the bilingual lexica (BILEXs) in the context of a number of languages including Sinhala. Yet another meta-analysis by Halpege [460] compared the commercially available English-Sinhala translation systems provided by Google and Bing through a comprehensive error analysis. This study used academic articles as its domain and ultimately concluded that both translation systems need significant improvements before they can be recommended for widespread usage.
@@
As mentioned in Section 3.1, the study by [99] raised questions on the quality of the existing Sinhala-English corpora. In a follow-up study by Ranathunga et al. [151], it was pointed out that using an automated ranking based on a similarity measure on web-mined corpora and using the resultant top samples can yield better translation models for English-Sinhala and Sinhala-Tamil. These models sometimes were better performing than a model trained on the full dataset. They further showed that using human labour to clean web-mined parallel corpora only gives marginal benefits over automated ranking and filtering. Thus they concluded that using expensive human labour for this task might not be efficient. The study by Team et al. [461] conducted a study on the quality of the translator models trained on their NLLB data set [130]. Naturally, this includes Sinhala.
Sen et al. [462] on the other hand attempted to improve the quality of the Sinhala-English parallel corpora using fuzzy string matching where they tried to match the English translation of the given Sinhala sentence to the English sentence in the dataset pair.
@@ 3.19.3 Singlish to Sinhala (Transliteration)
The XTREME-UP122122122https://github.com/google-research/xtreme-up data set created by Ruder et al. [137] contains a Sinhala data set for the transliteration task. The early work by Goonetilleke et al. [492] attempted Sinhala transliteration through the Latin alphabet. However, their work does not use the word transliteration and instead focuses on the predictive aspect. Priyadarshani et al. [493] used statistical machine learning for transliteration of names between Sinhala, Tamil, and English. A rule-based system on trigrams was proposed by Liwera and Ranathunga [494] for Singlish to Sinhala transliteration of social media text. A Singlish to Sinhala converter which uses an LSTM was proposed by De Silva [495]. A rule-based approach for the same was proposed by de Silva and Ahangama [496]. The study by Nanayakkara et al. [497] introduced an English-to-Sinhala transliteration system. Sumanathilaka et al. [498, 499] proposed a Trie [500] data structure-based algorithm for word suggestion in Sinhala transliterations named Swa-Bhasha. An extended analysis of the same work was presented in a later work [501]. A further extension, Swa-Bhasha 2.0, by Dharmasiri and Sumanathilaka [502] used NMT to eliminate word selection ambiguity in the transliteration process. Amarasekara et al. [503] proposed a rule-based method supported by N-gram analysis and a corpus dataset to transliterate Singlish tweets to Sinhala. The study by Rajapaksha et al. [504] claims to have created a translation system, however, what they have created is a transliteration system between Singlish and Sinhala. While they also claim to have trained ASR and TTS systems, the paper does not identify them by name or citation. The same is true for the data sets they have used to train the various modules in their pipeline. The studies by Kumaravithana et al. [505] and Jayawardhana et al. [506] claim to have built Sinhala-English translators but in actuality are just invoking Google API for their translation tasks.
The work by Kirov et al. [507] used the Dakshina [150] dataset and compared the accuracy of Singlish to Sinhala transliteration with vanilla LSTM, vanilla Transformers, mT5 [310], ByT5 [508], and a non-neural finite-state transducer (FST) based on work by Bisani and Ney [509]. They discuss in detail how the zero-width joiner (ZWJ, U+200D) character cause issues in this task and Sinhala rendering as a whole. Their code is available on GitHub123123123https://github.com/google-research/google-research/tree/master/context_aware_transliteration. Khiu et al. [510] fine-tuned mBART on Sinhala Government Corpus and Bible corpus to predict the NMT performance dependence on domain similarity.
@@ 3.19.5 Between Sinhala and Non-English Languages
Most of the cross-Sinhala and Tamil work has been done in the domain of machine translation. A neural machine translation for Sinhala and Tamil languages was initiated by Tennage et al. [514, 515]. Then they further enhanced it with transliteration and byte pair encoding [516] and used synthetic training data to handle the rare word problem [517]. This project produced Si-Ta [518] a machine translation system of Sinhala and Tamil official documents. In the statistical machine translation front, Farhath et al. [519] worked on integrating bilingual lists. The attempts by Weerasinghe [520] and Sripirakas et al. [521] were also focused on statistical machine translation while Jeyakaran and Weerasinghe [522] attempted a kernel regression method. A yet another attempt was made by Pushpananda et al. [523] which they later extended with some quality improvements [524]. An attempt at real-time direct translation between Sinhala and Tamil was done by Rajpirathap et al. [525]Dilshani et al. [526] have done a study on the linguistic divergence of Sinhala and Tamil languages with respect to machine translation. Mokanarangan [527] claims to have built a named entity translator between Sinhala and Tamil for official government documents. But this work is locked behind an institutional repository wall and thus is not accessible by other researchers. Arukgoda et al. [528] studied the possibility of using deep learning techniques to improve Sinhala-Tamil translation which they further improved later [529]Pramodya et al. [530] compared Transformers, Recurrent Neural Networks, and Statistical Machine Translation (SMT) in the context of Tamil to Sinhala machine translation. The work by Nissanka et al. [531] used monolingual word embedding to improve NMT between Sinhala and Tamil. Thillainathan et al. [532] uses pre-trained mBART [491] models for six directional translations between Sinhala, Tamil, and English. Yashothara and Uthayasanker [533] discussed the use of the Hierarchical Phrase-Based Model for Tamil to Sinhala and Sinhala to Tamil translations. Pramodya [534] presented a comparison of SMT [473] and NMT models for Sinhala-Tamil translation. For NMT they have used base transformer, turned transformer, and mT5 [310]. The work by Su et al. [535] which compared 8 Parameter-efficient fine-tuning (PEFT) methods on the Si-Ta language pair using the Government corpus [474] and NLLB [130] reports that Houlsby adapter [233], with a 33.34 SacreBLEU [536] score, to be the best PEFT method. However, they also note that the Pfeiffer adapter [235] runs the fastest at 52.59 hours with a reasonable score of 31.24. (Comparatively, the Houlsby adapter runs for 78.65 hours). The work by Pramodya et al. [537] conducted a comparative analysis of the available models for Sinhala-Tamil translation and reaffirmed the observations of Duh et al. [538] that, in low-resource scenarios, SMT and NMT both may work similarly, however, NMT needs more careful tuning to fit performance.
@@ 3.20 Spelling and Grammar
The open-source data-driven approach proposed by Wasala et al. [545, 546] claims to be able to check and correct spelling errors in Sinhala. The approach by Jayalatharachchi et al. [547] attempts to obtain synergy between two algorithms for the same purpose. These efforts [545, 547] were then extended by Subhagya et al. [548]. A rule-based Sinhala spell checker named SinSpell based on Hunspell124124124http://hunspell.github.io/ was introduced by Liyanapathirana et al. [549]. They have also made the tool available125125125http://nlp-tools.uom.lk/sinspell/ online for use. The study by Sithamparanathan and Uthayasanker [550] extended the Generic Environment for context-aware spell correction to handle Sinhala and Tamil. Sonnadara et al. [123] created a benchmark data set for Sinhala spell correction along with a neural model. A multi agent-based spell checker, named LaSi Spell for Sinhala spell checking was introduced by Samarawickrama et al. [551]. The study by Udagedara et al. [552] specifically solved the problem of spell-checking Sri Lankan names and addresses. A system named Erroff was proposed by Sudesh et al. [553] to correct real-word errors in Sinhala text.
A model for detecting grammatical mistakes in Sinhala was developed by Pabasara and Jayalal [554]. They followed this up with a grammatical error detection and correction model [555]Gunasekara et al. [556] used annotation projection for semantic role labelling for Sinhala. A Sinhala grammar checker based on Hidden Markov models was developed by Fernando and Arudchelvam [557]. Widyaratna [558] used a sequence-to-sequence model with attention, which is generally used for translation tasks, to translate sentences with common grammatical errors to their corrected counterparts. The work by Jayasuriya et al. [559] used a rule-based approach along with Google Translation to correct grammar in Sinhala text. A rule-based system to convert Sinhala sentences from active voice to passive voice while correcting grammatical errors was proposed by Ilukkumbura and Rupasinghe [560]. The study by Goonawardena et al. [561] a rule-based system to spell-check Sinhala text as well as detect and correct grammatical errors. Navoda et al. [562] also claims to have created an automated tool to check Sinhala spelling and grammar.
@@ 3.21 Chat Bots
A simple Sinhala chat bot which utilizes a small knowledge base has been proposed by Hettige and Karunananda [73]. A study on the effect of word embeddings on a Sinhala chatbot was conducted by Gamage et al. [563] where they used, the fasttext model trained by Facebook [116, 117, 118], on a RASA126126126https://rasa.com/ [564] chat bot. A Sinhala chat bot for train information was proposed by Harshani [565]. Similarly, the tool proposed by Chandrasena et al. [566] serves as a chat bot-based recommendation system for Sri Lankan traditional dancers. The chat bot discussed by Kumanayake [567] has the very specific purpose of answering user inquiries about the degree programs at University of Ruhuna. Avishka et al. [568] used off the shelf RASA NLU Engine [564] and Microsoft Bot Network [569] to set up a generic Sinhala chat bot architecture. They demonstrated the effectiveness of their architecture by creating a food ordering chat bot. A web-based code-less chat bot development platform for Sinhala was proposed by Dissanayake et al. [570]. Further, they claimed that their system can handle OOV tokens as well as Sinhala-English code-switching. The work by Dasanayaka and Warnajith [571] used a deep learning Intent Mapping (IM) model to map consumer responses in their Sinhala chat bot framework. Rajapakshe et al. [375] proposed a Sinhala conversational interface for the purpose of scheduling medical appointments and giving medical advice. The chat bot component was in the back end while the front end contained speech recognition and voice synthesizer components.